47. Calculate the rotational inertia (kg* m^2) of a wheel that has kinetic energy of 15,000J when rotating at 502 rpm. (in kg*m^2)
Answer #1

angular speed, w = 502 rpm

= 502*2*pi/60 rad/s

= 52.54 rad/s

we know, Rotational Kinetic enrgy, KE = 0.5*I*w^2

===> I = 2*KE/w^2

= 2*15000/52.54^2

= 10.87 kg.m^2 <<<<<<<<<-----------Answer